OpenStudy (dls):

A stationary pulley carries a rope,one end of which supports a ladder with a man and the other a counter weight mass M.The man of mass 'm' climbs up a distance "L" wrt the ladder and then stops.The displacement of CM of the system is?

OpenStudy (dls):

@Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy @Mashy

OpenStudy (anonymous):

yabba.. do it man.. first draw it!

OpenStudy (dls):

Options: a ) \[\frac{ml}{M+2m}\] b)\[\frac{ml}{2M}\] c)\[\frac{ml}{M+m} \] d) \[\frac{ml}{2M+m}\]

OpenStudy (dls):

I cant even draw the diagram!!

OpenStudy (dls):

is it in air?

OpenStudy (anonymous):

yea.. put the pulley in the air

OpenStudy (anonymous):

wait.. ll draw and put photo :P

OpenStudy (dls):

WTH I MADE AN AWESOME DIAGRAM AND pressed back -_-

OpenStudy (dls):

GTH

OpenStudy (anonymous):

ok uploading wait!

OpenStudy (dls):